# Compact Kähler manifolds with quasi-positive holomorphic sectional curvature

### v3

@@ 2.3. Some properties of curvature tensor
$\partial_{u}\partial_{\bar{u}}|\psi|^{2}$
$=$ $\langle D^{\prime}_{u}\psi,\overline{D^{\prime}_{u}\psi}\rangle+\sum\limits_{i,j=1}^{n}R^{h}_{u\bar{u}i\bar{j}}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,j),J_{l}}}$
@@
with $\iota^{p+k}_{(e^{t})^{*}}(\psi)=\psi(\cdots;\cdots,(e^{t})^{*},\cdots)$ where the contractions are done at the $(p+k)$-th slot. Now we can reselect the local holomorphic coordinates of $M$ and the local holomorphic basis of $E$ such that $(h_{i\bar{j}}),(H_{t\bar{s}})$ are still the identity and $(\alpha_{i\bar{j}}),(\beta_{t\bar{s}})$ are diagonal at $x$, then
$\sum\limits_{i=1}^{n}\alpha_{i\bar{i}}=\sum\limits_{i=1}^{n}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,i),J_{l}}}=p|\psi|^{2}$ (5.5)
and
$\sum\limits_{t=1}^{r}\beta_{t\bar{t}}=\sum\limits_{t=1}^{r}\sum\limits_{k=1}^{l}\sum\limits_{I_{p},J_{l-1}}a_{I_{p},J_{l-1}(k,t)}\overline{a_{I_{p},J_{l-1}(k,t)}}=l|\psi|^{2}.$ (5.6)
So, for any holomorphic section $\psi\in H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})$,

### v4

@@ 2.4. The structure theorem of morphism
$\partial_{u}\partial_{\bar{u}}|\psi|^{2}=$ $\langle D^{\prime}_{u}\psi,\overline{D^{\prime}_{u}\psi}\rangle+\sum\limits_{i,j=1}^{n}R^{h}_{u\bar{u}i\bar{j}}\sum\limits_{m=1}^{p}\sum\limits_{I_{p-1},J_{l}}a_{I_{p-1}(m,i),J_{l}}\overline{a_{I_{p-1}(m,j),J_{l}}}$
@@
with $\iota^{p+k}_{(e^{t})^{*}}(\psi)=\psi(\cdots;\cdots,(e^{t})^{*},\cdots)$ where the contractions are done at the $(p+k)$-th slot. Now we can reselect the local holomorphic coordinates of $M$ and the local holomorphic basis of $E$ such that $(h_{i\bar{j}}),(H_{t\bar{s}})$ are still the identity and $(\alpha_{i\bar{j}}),(\beta_{t\bar{s}})$ are diagonal at $x$, then $\sum\limits_{i=1}^{n}\alpha_{i\bar{i}}=p|\psi|^{2}$ and $\sum\limits_{t=1}^{r}\beta_{t\bar{t}}=l|\psi|^{2}.$ So, for any holomorphic section $\psi\in H^{0}(M,((T^{1,0}M)^{*})^{\otimes p}\otimes E^{\otimes l})$,
@@
Appendix. Proof of $L^{2}$-vanishing Theorem
Proof of Proposition 4.1.
Since $\omega$ is complete, there exists an exhaustive sequence $(K_{v})_{v\in\mathbb{N}}$ of compact subsets of $M$ and functions $\psi_{v}\in C^{\infty}(M,\mathbb{R})$ (see. e.g. [14, Section 8.2]) such that
  • $\psi_{v}=1$ in a neighborhood of $K_{v}$, $\mathrm{supp}\psi_{v}\subset K^{\circ}_{v+1}$,
  • $0\leq\psi_{v}\leq 1$ and $|d\psi_{v}|^{2}_{g}\leq 2^{-v}\psi_{v}$ on $M$.
Let $\eta$ be a nonzero $L^{2}$-holomorphic $(p,0)$-form, then $|\eta|\in W^{1,\infty}_{loc}(M)$ and $\Sigma:=\{x\in M:\eta(x)=0\}$ is a proper subvariety of $M$, from Lemma 3.2 and Bochner-type formula (3.14), we have
$\begin{split}&-\int_{M}\langle\partial\psi_{v},\bar{\partial}|\eta|_{g}^{2}\rangle_{g}\cdot\omega^{n}=\int_{M}\psi_{v}\Delta|\eta|_{g}^{2}\cdot\omega^{n}\\ \geq&\int_{M}\psi_{v}\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\frac{\beta}{|\eta|_{g}^{2}}\rangle\cdot\omega^{n}+\int_{M}\psi_{v}\cdot\kappa|\eta|_{g}^{2}\cdot\omega^{n}+\frac{n!(\sqrt{-1})^{p^{2}}}{(n-p-1)!}\int_{M}\psi_{v}\sqrt{-1}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}\\ \geq&\int_{M}\psi_{v}\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\frac{\beta}{|\eta|_{g}^{2}}\rangle\cdot\omega^{n}+\frac{n!(\sqrt{-1})^{p^{2}}}{(n-p-1)!}\int_{M}\psi_{v}\sqrt{-1}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}\end{split}$ (5.6)
where $\beta=\Lambda^{p-1}(\frac{\eta\wedge\overline{\eta}}{p!})$. Recall (3.3) that $\operatorname{tr}_{\omega}\beta=|\eta_{g}|^{2}$, then $|\frac{\beta}{|\eta|_{g}^{2}}|_{g}\leq 1$ and $|\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}|_{g}\leq 1$ on $M\setminus\Sigma$, thus each integral on the RHS of (5.6) makes senses. For any fixed point $x_{0}\in M$, we choose local holomorphic coordinates $(z^{1},\cdots,z^{n})$ centered at $x_{0}$ such that $(g_{i\bar{j}})$ is the identity and $(\beta_{i\bar{j}})$ is diagonal, then
$4\langle\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g},\beta\rangle_{g}=\frac{\partial_{i}|\eta|_{g}^{2}\cdot{\partial}_{\bar{i}}|\eta|_{g}^{2}}{|\eta|_{g}^{2}}\beta_{i\bar{i}}=\frac{\langle D^{\prime}_{i}\eta,\bar{\eta}\rangle\cdot\langle\eta,\overline{D^{\prime}_{i}\eta}\rangle}{|\eta|_{g}^{2}}\beta_{i\bar{i}}\leq\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\beta\rangle.$ (5.7)
As the $L^{2}$-holomorphic $(p,0)$-form $\eta$ is closed (see [20]), applying Lebesgue’s dominated covergence theorem, we have
$\begin{split}0=&\int_{M}\partial(\psi_{v}\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1})\\ =&\lim_{\epsilon\rightarrow 0}\int_{M}\partial(\psi_{v}\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}+\epsilon}\wedge\omega^{n-p-1})\\ =&\lim_{\epsilon\rightarrow 0}\int_{M}\left(\partial\psi_{v}\wedge\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}+\epsilon}+\psi_{v}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}+\epsilon}-4\psi_{v}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g}\wedge\frac{|\eta|_{g}^{2}}{(|\eta|_{g}^{2}+\epsilon)^{2}}\eta\wedge\bar{\eta}\right)\wedge\omega^{n-p-1}\\ =&\int_{M}\left(\partial\psi_{v}\wedge\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}+\psi_{v}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}-4\psi_{v}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\right)\wedge\omega^{n-p-1}.\end{split}$ (5.8)
Applying Lemma 3.2, we have
$\int_{M}\psi_{v}|\partial|\eta|_{g}|_{g}^{2}\cdot\omega^{n}=\int_{M}\psi_{v}\operatorname{tr}_{\omega}(\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g})\cdot\omega^{n}$
$=$ $\int_{M}\psi_{v}\langle\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g},\frac{\beta}{|\eta|_{g}^{2}}\rangle_{g}\cdot\omega^{n}+\frac{(\sqrt{-1})^{p^{2}}n!}{(n-p-1)!}\int_{M}\psi_{v}\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}$
$\leq$ $\frac{1}{4}\left(\int_{M}\psi_{v}\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\frac{\beta}{|\eta|_{g}^{2}}\rangle\cdot\omega^{n}+\frac{(\sqrt{-1})^{p^{2}}n!}{(n-p-1)!}\lim_{\epsilon\rightarrow 0}\int_{M}\psi_{v}\sqrt{-1}\partial|\eta|_{g}\wedge\bar{\partial}|\eta|_{g}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}+\epsilon}\wedge\omega^{n-p-1}\right)$
$=$ $\frac{1}{4}\left(\int_{M}\psi_{v}\langle\sqrt{-1}\langle D^{\prime}\eta,\overline{D^{\prime}\eta}\rangle,\frac{\beta}{|\eta|_{g}^{2}}\rangle\cdot\omega^{n}+\frac{n!(\sqrt{-1})^{p^{2}}}{(n-p-1)!}\int_{M}\psi_{v}\sqrt{-1}\partial\bar{\partial}|\eta|_{g}^{2}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}\right)$
$-\frac{(\sqrt{-1})^{p^{2}}n!}{4(n-p-1)!}\int_{M}\sqrt{-1}\partial\psi_{v}\wedge\bar{\partial}|\eta|^{2}_{g}\wedge\frac{\eta\wedge\bar{\eta}}{|\eta|_{g}^{2}}\wedge\omega^{n-p-1}$
$\leq$ $C_{1}\int_{M}|\partial\psi_{v}|_{g}\cdot\big{|}\partial|\eta|_{g}\big{|}\cdot|\eta|_{g}\cdot\omega^{n}$
for some constant $C_{1}>0$ independent of $v\in\mathbb{N}$. The fact that $|\partial\psi_{v}|_{g}^{2}\leq 2^{-v}\psi_{v}$ and the Schwarz inequality yields
$\int_{M}\psi_{v}\big{|}\partial|\eta|_{g}\big{|}_{g}^{2}\cdot\omega^{n}\leq 2^{-v}C_{1}||\eta||_{L^{2}},$
hence $\int_{M}\psi_{v}\big{|}\partial|\eta|_{g}\big{|}_{g}^{2}\cdot\omega^{n}\rightarrow 0$ for $v\rightarrow\infty$, which implies that $|\eta|_{g}\equiv C$ for some constant $C\geq 0$. Applying the pointwise computation in the proof of Theorem 1.4, the fact (3.25) tells us that $D\eta=0$. The proof is completed. ∎